Solve each system by elimination. Check your answer.
3x+3y=15 and -2x+3y=-5

Answer by thair(3)

by multipling eq 1 by 2 and eq. 2 by 3 we get
6x+6y=30
-6x+9y=-15 and adding both of them to remove x we got 15y=15 , y=1 substut this value in any eq.let use eq 1 3x+1*3=15 ,x=(15-3)/3=4
Ans y=1,x=4